OpenStudy (anonymous):

suppose f(x)=4x-2 and g(x)=-2x+1 Find the value of f(5)/g(-3) A. 1 5/9 B. 2 4/7 C. -2 D. 2

OpenStudy (anonymous):

its simple just put the values in and solve

OpenStudy (mayankdevnani):

@jurneelynnn given eq. is f(x)=4x-2 then put x=5 then we get, f(5)=4*5-2=???? can you solve it???

OpenStudy (mayankdevnani):